### v1

@@ Abstract
In this paper, we address the issue of estimating and inferring the distributional treatment effects in randomized experiments. The distributional treatment effect provides a more comprehensive understanding of treatment effects by characterizing heterogeneous effects across individual units, as opposed to relying solely on the average treatment effect. To enhance the precision of distributional treatment effect estimation, we propose a regression adjustment method that utilizes the distributional regression and pre-treatment information. Our method is designed to be free from restrictive distributional assumptions. We establish theoretical efficiency gains and develop a practical, statistically sound inferential framework. Through extensive simulation studies and empirical applications, we illustrate the substantial advantages of our method, equipping researchers with a powerful tool for capturing the full spectrum of treatment effects in experimental research.
@@ 1 Introduction
Our paper provides several important contributions to the existing literature. First, our paper contributes to a growing body of literature on heterogeneous treatment effects in econometrics and statistics (e.g., Athey and Imbens, 2016; Wager and Athey, 2018). The quantile treatment effect (QTE) was first introduced by Doksum (1974) and Lehmann (1975) for estimating treatment effects due to unobserved heterogeneity. Since then, various estimation and inference methods for the distributional and quantile treatment effect have been developed and applied, including Heckman et al. (1997), Abadie et al. (2002), Athey and Imbens (2006), Bitler et al. (2006), Chernozhukov and Hansen (2005), Djebbari and Smith (2008), Donald and Hsu (2014), Callaway et al. (2018), Callaway and Li (2019), Chernozhukov et al. (2020), and Kallus et al. (2024), among others. Jiang et al. (2023) studies the estimation of regression-adjusted QTE under covariate adaptive randomization schemes. To the best of our knowledge, however, techniques for estimating DTE has not been explored in the context of regression-adjustment method.666 To estimate heterogeneous treatment effects, researchers commonly use two mututally not exclusive approaches: conditional average treatment effect (CATE) and distributional treatment effect (DTE). CATE estimates the average treatment effect for subgroups based on observed variables (Athey and Imbens, 2016; Imai and Ratkovic, 2013; Künzel et al., 2019; Nie and Wager, 2020; Wager and Athey, 2018). DTE, on the other hand, compares outcome distributions between treatment and control groups across the entire spectrum, and thus it can capture effects that CATE might miss, particularly those stemming from unobserved factors. As such, DTE remains valuable even after accounting for observed variables in CATE analyses.
@@ 2.1 Setup and Treatment Effects
where the functional $\Psi_{k,k^{\prime}}^{DTE}:\Gamma\to\ell^{\infty}(\mathcal{Y})$ is defined as $\Psi_{k,k^{\prime}}^{DTE}(\gamma)=F_{Y(k)}-F_{Y(k^{\prime})}$. One of its advantages is that DTE is well-defined for any type of outcome, including discrete variables.777 Our approach measures the distributional effect as the difference between the distributions of potential outcomes under distinct treatments. While this paper focuses on DTE, we note that under the rank invariance assumption, our results are also informative about the distribution of individual treatment effects (ITE). Without additional conditions or rank invariance, researchers may consider adopting a partial-identification approach for estimating the distribution of ITE. See Heckman et al. (1997) among others.
@@
PTE measures changes in the probability that the outcome variable falls in interval $(y,y+h]$ and can apply for any type of outcome variables. For an integer outcomes and $h=1$, $\Delta_{k,k^{\prime}}^{PTE}(y,1)$ quantifies the probability change at location $y$, or $\Pr\{Y(k)=y\}-\Pr\{Y(k^{\prime})=y\}$. Alternatively, when the outcome variable is continuous, we can consider QTE, defined through a simple inversion of the distribution function: for $u\in(0,1)$, $\Delta^{QTE}_{k,k^{\prime}}(u):=F_{Y(k)}^{-1}(u)-F_{Y(k^{\prime})}^{-1}(u),$ where $u\mapsto F_{Y(k)}^{-1}(u)$ is the quantile function of $Y(k)$. Since standard analysis based on quantile functions presumes the existence of the outcome density, QTE is not readily applicable in the presence of discreteness in the outcome distributions.
@@ 2.2 Distributional Regression
In many practical applications, the true conditional distribution $F_{Y(k)|\bm{X}}(y|\bm{X})$ remains unknown. It is crucial not to impose overly restrictive global parametric constraints. To address this issue, we employ the distribution regression framework, wherein a parametric linear-index model is used to describe the binary outcomes $1\!{\rm l}_{\{Y(k)\leq y\}}$ for each $y\in\mathcal{Y}$. More specifically, letting $\Lambda:\mathbb{R}\to[0,1]$ be a known link function, we define the following conditional distribution based on a linear index model for each fixed $y\in\mathcal{Y}$:
@@
where $T:\mathcal{X}\to\mathbb{R}^{d}$ is some transformation, such as polynomial or pair-wise interaction of regressors, specified by researchers, and $\beta_{k}(y){\in}\mathbb{R}^{d}$ are unknown parameters. Setting the link function as the normal or logistic distribution function, we can consider the models as probit or logit models for binary outcome $1\!{\rm l}_{\{Y(k)\leq y\}}$, respectively.
@@
where ^$\widehat{\ell}_{k}(\beta;y):=n_{k}^{-1}\sum_{i=1}^{n}W_{k,i}\cdot\ell_{i}(\beta;y)$ is the log-likelihood function for treatment $k\in\mathcal{K}$, and $\mathcal{B}\subset\mathbb{R}^{d}$ is the parameter space. Given the estimator $\widehat{\beta}_{k}(y)$, we define the conditional distribution function estimator as
@@
We can estimate the conditional distribution above for each value of $y\in\mathcal{Y}$ when the outcome variable $Y$ takes finite discrete values, and we select a sufficiently large set of discrete points $\{y_{j}\in\mathcal{Y}\}_{j=1}^{J}$ for estimation when the outcome variable is a continuous random variable.888 One important property of the conditional distribution function is non-decreasing by definition, while the estimator $\widehat{G}_{Y(k)|\bm{X}}(y|\bm{x})$ does not necessarily satisfy monotonicity in finite samples due to estimation errors. We can monotonize the conditional distribution estimators using the rearrangement method, for example, as proposed by Chernozhukov et al. (2009).
@@ 2.3 Regression-Adjusted Treatment Effect Estimator
^$\widehat{\mathbb{P}}_{\bm{X}}f=\frac{1}{n}\sum_{i=1}^{n}f(\bm{X}_{i})\ \ \ \ \mathrm{and}\ \ \ \ \widehat{\mathbb{P}}_{\bm{X}}^{(k)}f=\frac{1}{n_{w}}\sum_{i=1}^{n}W_{k,i}\cdot f(\bm{X}_{i}).$
@@
The linear index model in (3) can be considered as one of the generalized linear models (GLM). In the literature, it is well known that the certain choice of specifications, such as a link function, can guarantee the unbiased estimator for the unconditional mean in finite samples. Such choice of the link function is called the canonical link function (see McCullagh and Nelder, 1989). More precisely, the canonical link function should yield
@@
At the population level, this relation can be expressed as: $\mathbb{E}\big{[}1\!{\rm l}_{\{Y(k)\leq y\}}-G_{Y(k)|\bm{X}}(y|\bm{X})\big{]}=0.$ This property is important for the regression adjustment to guarantee that the regression adjustment does not induce bias to improve the precision of the treatment effect estimator. Negi and Wooldridge (2021) pointed this out for a nonlinear regression adjustment method and derive the asymptotic property when the true distribution is known. In what follows, we focus on the canonical class of the GLM model for binary outcome, which includes OLS and logit, while we do not assume any global parametric assumptions and allow for model mis-specification.
@@ 3.1 Oracle Estimator and Efficiency
where the equality holds only if $G_{Y(k)|\bm{X}}=F_{Y(k)}$.
@@
Theorem 1(a) shows the efficiency gains achieved by applying regression adjustment to estimate unconditional distribution functions. More specifically, the variance of the unconditional distribution estimators decreases if $G_{Y(k)|\bm{X}}(y)$ has some predictive power for $\mathbbm{1}_{\{Y(k)\leq y\}}$. Furthermore, Theorem 1(b) elaborates on these gains in a positive semi-definite sense for a vector of regression-adjusted estimators. This result encompasses any linear combination of the unconditional distribution estimators and implies potential efficiency gains for both DTE and PTE estimators. While a similar positive-definiteness result for non-linear regression adjustment is reported in Theorem 6 of Negi and Wooldridge (2020), our result provides an additional insight that the efficiency gains occur when the regression models are not linearly dependent across treatments. Theorem 1(b) requires correct specification, which is analogous to the requirement in Negi and Wooldridge (2020) for non-linear regression adjustment.
@@ 3.2 Asymptotic Distribution of DTE Estimator
(a) The log-likelihood functions $\beta\mapsto\ell_{k}(\beta;y)$ is a concave function. The link function $\Lambda(\cdot)$ belongs to the canonical class and is twice continuously differentiable with its first derivative $\lambda(\cdot)$.
@@
Suppose that Assumptions 1 and 2 hold. Then, we have, in $\ell^{\infty}(\mathcal{Y})$,
$\sqrt{n}(\widehat{\Delta}^{DTE}_{k,k^{\prime}}-\Delta^{DTE}_{k,k^{\prime}})\rightsquigarrow\Psi^{DTE}_{k,k^{\prime}}(\mathbb{B}),$
$\sqrt{n}(\widehat{\Delta}^{PTE}_{k,k^{\prime},h}-\Delta^{PTE}_{k,k^{\prime},h})\rightsquigarrow\Psi^{PTE}_{k,k^{\prime},h}(\mathbb{B}),$
for every $k,k^{\prime}\in\mathcal{K}$ and some constant $h>0$, where $\mathbb{B}$ is a mean-zero Gaussian process defined in Lemma A.4 of Supplemental Material.
@@
Let $(S_{1},\dots,S_{n})$ be $n$ scalar, nonnegative random variables, which are independent and identically distributed and also independent of the original sample. The random weights $\{S_{i}\}_{i=1}^{n}$ satisfy the followings: $\mathbb{E}|S_{i}|^{2+\epsilon}<\infty$ for some $\epsilon>0$, $\bar{S}_{n}:=n^{-1}\sum_{i=1}^{n}S_{i}\to^{p}0$ and $n^{-1}\sum_{i=1}^{n}(S_{i}-\bar{S}_{n})^{2}\to^{p}1$.
@@
In the above equation, we use $\widehat{G}_{Y(k)|\bm{X}}^{\ast}$ the conditional distribution estimator based on the original sample, instead of the bootstrap sample. According to Theorem 2, the conditional distribution estimator has no impact on the limit distribution and we keep the estimator throughout the resampling procedure. The validity is proven in the theorem below.
@@
$\widehat{\Delta}^{DTE\ast}_{k,k^{\prime}}:=\Psi_{k,k^{\prime}}^{DTE}(\,\widehat{\gamma}^{\ast}\,)\ \ \ \mathrm{and}\ \ \ \widehat{\Delta}^{PTE\ast}_{k,k^{\prime},h}:=\Psi_{k,k^{\prime},h}^{PTE}(\,\widehat{\gamma}^{\ast}\,).$
In the theorem provided below, we show that the exchangeable bootstrap provides a method to consistently estimate the limit process of the regression-adjusted DTE and PTE estimators, following van der Vaart and Wellner (1996).
@@
Under Assumptions 1-3, we have, in $\ell^{\infty}(\mathcal{Y})$,
@@
[…]
Theorems 2 and 3 present a methodologically rigorous and empirically useful approach for estimating DTE and PTE. Moreover, these theoretical results establish the asymptotic validity of the exchangeable bootstrap procedure for constructing confidence intervals and conducting hypothesis tests for both DTE and PTE.

### v2

@@ Abstract
In this paper, we address the issue of estimating and inferring distributional treatment effects in randomized experiments. The distributional treatment effect provides a more comprehensive understanding of treatment heterogeneity compared to average treatment effects. We propose a regression adjustment method that utilizes distributional regression and pre-treatment information, establishing theoretical efficiency gains without imposing restrictive distributional assumptions. We develop a practical inferential framework and demonstrate its advantages through extensive simulations. Analyzing water conservation policies, our method reveals that behavioral nudges systematically shift consumption from high to moderate levels. Examining health insurance coverage, we show the treatment reduces the probability of zero doctor visits by 6.6 percentage points while increasing the likelihood of 3-6 visits. In both applications, our regression adjustment method substantially improves precision and identifies treatment effects that were statistically insignificant under conventional approaches.
@@ 1 Introduction
Our paper provides several important contributions to the existing literature. First, our paper contributes to a growing body of literature on heterogeneous treatment effects in econometrics and statistics (e.g., Athey and Imbens, 2016; Wager and Athey, 2018). The quantile treatment effect (QTE) was first introduced by Doksum (1974) and Lehmann (1975) for estimating treatment effects due to unobserved heterogeneity. Since then, various estimation and inference methods for the distributional and quantile treatment effect have been developed and applied, including Heckman et al. (1997), Abadie et al. (2002), Athey and Imbens (2006), Bitler et al. (2006), Chernozhukov and Hansen (2005), Djebbari and Smith (2008), Donald and Hsu (2014), Callaway et al. (2018), Callaway and Li (2019), Chernozhukov et al. (2020), and Kallus et al. (2024), among others. Jiang et al. (2023) study the estimation of regression-adjusted QTE under covariate adaptive randomization schemes. To the best of our knowledge, however, techniques for estimating DTE has not been explored in the context of regression-adjustment method.666 To estimate heterogeneous treatment effects, researchers commonly use two mututally not exclusive approaches: conditional average treatment effect (CATE) and distributional treatment effect (DTE). CATE estimates the average treatment effect for subgroups based on observed variables (Athey and Imbens, 2016; Imai and Ratkovic, 2013; Künzel et al., 2019; Nie and Wager, 2020; Wager and Athey, 2018). DTE, on the other hand, compares outcome distributions between treatment and control groups across the entire spectrum, and thus it can capture effects that CATE might miss, particularly those stemming from unobserved factors. As such, DTE remains valuable even after accounting for observed variables in CATE analyses. , [^7] Jiang et al. (2023) and our paper complement each other. While Jiang et al. (2023) considers our experimental scheme as a special case and employs doubly robust moment conditions for quantile regression, our proposed method applies to non-continuous outcome variables and also employs doubly robust moment condition with distribution regression framework.
@@ 2.1 Setup and Treatment Effects
where the functional $\Psi_{k,k^{\prime}}^{DTE}:\Gamma\to\ell^{\infty}(\mathcal{Y})$ is defined as $\Psi_{k,k^{\prime}}^{DTE}(\gamma)=F_{Y(k)}-F_{Y(k^{\prime})}$. One of its advantages is that DTE is well-defined for any type of outcome, including discrete variables.888 Our approach measures the distributional effect as the difference between the distributions of potential outcomes under distinct treatments. While this paper focuses on DTE, we note that under the rank invariance assumption, our results are also informative about the distribution of individual treatment effects (ITE). Without additional conditions or rank invariance, researchers may consider adopting a partial-identification approach for estimating the distribution of ITE. See Heckman et al. (1997) among others.
@@
PTE measures changes in the probability that the outcome variable falls in interval $(y,y+h]$ and can apply for any type of outcome variables. For an integer outcomes and $h=1$, $\Delta_{k,k^{\prime},1}^{PTE}(y)$ quantifies the probability change at location $y$, or $\Pr\{Y(k)=y\}-\Pr\{Y(k^{\prime})=y\}$. Alternatively, for continuous outcome variables, we can consider the quantile treatment effect (QTE), which is defined through a simple inversion of the distribution function. For $u\in(0,1)$, let the quantile function of $Y(k)$ be defined as the map $u\mapsto F_{Y(k)}^{-1}(u)$. The QTE between treatments $k,k^{\prime}\in\mathcal{K}$ is defined as $\Delta^{QTE}_{k,k^{\prime}}:=\Psi_{k,k^{\prime}}^{QTE}(\gamma)(u)$ where $\Psi_{k,k^{\prime}}^{QTE}(\gamma)(u):=F_{Y(k)}^{-1}(u)-F_{Y(k^{\prime})}^{-1}(u).$ Since standard analysis based on quantile functions presumes the existence of the outcome density, QTE is not readily applicable in the presence of discreteness in the outcome distributions.
@@ 2.2 Distributional Regression
In many practical applications, the true conditional distribution $F_{Y(k)|\bm{X}}(y|\bm{X})$ remains unknown. It is crucial not to impose overly restrictive global parametric constraints. To address this issue, we employ the distribution regression framework, wherein a parametric linear-index model is used to describe the binary outcomes $1\!{\rm l}_{\{Y(k)\leq y\}}$ for each $y\in\mathcal{Y}$. More specifically, let $\Lambda:\mathbb{R}\to[0,1]$ be a known inverse link function. For each $y\in\mathcal{Y}$, we define the following conditional distribution based on a linear index model:
@@
where $T:\mathcal{X}\to\mathbb{R}^{d}$ is some transformation, such as polynomial or pair-wise interaction of regressors, specified by researchers, and $\beta_{k}(y){\in}\mathbb{R}^{d}$ are unknown parameters. Setting $\Lambda(\cdot)$ as either the normal or logistic distribution function yeilds probit or logit models, respectively, for binary outcome $1\!{\rm l}_{\{Y(k)\leq y\}}$.
@@
where the log-likelihood function for treatment $k$ is denoted by ^$\widehat{\ell}_{k}(\beta;y):=n_{k}^{-1}\sum_{i=1}^{n}W_{k,i}\cdot\ell_{i}(\beta;y)$ with $\ell_{i}(\beta:y)=1\!{\rm l}_{\{Y_{i}\leq y\}}\log\Lambda\big{(}T(\bm{X}_{i})^{\top}\beta\big{)}+(1-1\!{\rm l}_{\{Y_{i}\leq y\}})\log\big{\{}1-\Lambda\big{(}T(\bm{X}_{i})^{\top}\beta\big{)}\big{\}}$ and $\mathcal{B}\subset\mathbb{R}^{d}$ is the parameter space. Given the estimator $\widehat{\beta}_{k}(y)$, we define the conditional distribution function estimator as
@@
We can estimate the conditional distribution above for each value of $y\in\mathcal{Y}$ when the outcome variable $Y$ takes finite discrete values, and we select a sufficiently large set of discrete points $\{y_{j}\in\mathcal{Y}\}_{j=1}^{J}$ for estimation when the outcome variable is a continuous random variable.999 One important property of the conditional distribution function is non-decreasing by definition, while the estimator $\widehat{G}_{Y(k)|\bm{X}}(y|\bm{x})$ does not necessarily satisfy monotonicity in finite samples due to estimation errors. We can monotonize the conditional distribution estimators using the rearrangement method, for example, as proposed by Chernozhukov et al. (2009).
@@ 2.3 Regression-Adjusted Treatment Effect Estimator
^$\widehat{\mathbb{P}}_{\bm{X}}f=\frac{1}{n}\sum_{i=1}^{n}f(\bm{X}_{i})\ \ \ \ \mathrm{and}\ \ \ \ \widehat{\mathbb{P}}_{\bm{X}}^{(k)}f=\frac{1}{n_{k}}\sum_{i=1}^{n}W_{k,i}\cdot f(\bm{X}_{i}).$
@@
Similarly, the QTE estimator is defined as $\widehat{\Delta}^{QTE}_{k,k^{\prime}}:=\Psi_{k,k^{\prime}}^{QTE}(\,\widehat{\gamma}\,)$.
The linear index model in (3) can be viewed as a special case of generalized linear models (GLMs). Within the GLM framework, using the canonical link function yields unbiased estimation of the unconditional mean in finite samples (see Section A.5 for more details and McCullagh and Nelder, 1989). Specifically, the canonical link function must satisfy
@@
At the population level, this relation can be expressed as: $\mathbb{E}\big{[}1\!{\rm l}_{\{Y(k)\leq y\}}-G_{Y(k)|\bm{X}}(y|\bm{X})\big{]}=0.$ This property is important for the regression adjustment to guarantee that the regression adjustment does not induce bias to improve the precision of the treatment effect estimator. Negi and Wooldridge (2021) pointed this out for a nonlinear regression adjustment method and derive the asymptotic property when the true distribution is known. In what follows, we focus the GLMs with the canonical link functions for binary outcome, which includes OLS and logit, while we do not assume any global parametric assumptions and allow for model mis-specification.
@@ 3.1 Oracle Estimator and Efficiency
provided that $\mathbb{E}[(1\!{\rm l}_{\{Y(k)\leq\cdot\}}-F_{Y(k)})^{2}]\geq\mathbb{E}[(1\!{\rm l}_{\{Y(k)\leq\cdot\}}-G_{Y(k)|\bm{X}})^{2}]$.
@@
Theorem 1(a) shows the efficiency gains achieved by applying regression adjustment to estimate unconditional distribution functions. More specifically, the variance of the unconditional distribution estimators decreases if $G_{Y(k)|\bm{X}}(y)$ has some predictive power for $1\!{\rm l}_{\{Y(k)\leq y\}}$ in the $L^{2}$ sense. When $\Lambda(\cdot)$ is the identity function or in the case of a linear prediction model, this property holds automatically. In the case of logistic regression, the level of model misspecification must be sufficiently small relative to the predictive signal in the data.101010 A simple algebra shows that $\mathbb{E}[(1\!{\rm l}_{\{Y(k)\leq\cdot\}}-F_{Y(k)})^{2}]-\mathbb{E}[(1\!{\rm l}_{\{Y(k)\leq\cdot\}}-G_{Y(k)|\bm{X}})^{2}]=\mathbb{E}[(F_{Y(k)|\bm{X}}-F_{Y(k)})^{2}]-\mathbb{E}[(F_{Y(k)|\bm{X}}-G_{Y(k)|\bm{X}})^{2}]$. Here, the first term measures the signal in $\bm{X}$ for predicting $1\!{\rm l}_{\{Y(k)\leq\cdot\}}$. The second term quantifies the model misspecification, that is, the $L^{2}$ distance between the true conditional distribution $F_{Y(k)|\bm{X}}$ and the specified model $G_{Y(k)|\bm{X}}$. When the model misspecification (second term) is sufficiently small relative to the signal in the data (quantified by the variance of $F_{Y(k)|\bm{X}}$), the predictor $G_{Y(k)|\bm{X}}$ achieves variance reduction for the unconditional estimator.
Furthermore, Theorem 1(b) elaborates on these gains in a positive semi-definite sense for a vector of regression-adjusted estimators. This result encompasses any linear combination of the unconditional distribution estimators and implies potential efficiency gains for both DTE and PTE estimators. While a similar positive-definiteness result for non-linear regression adjustment is reported in Theorem 6 of Negi and Wooldridge (2020), our result provides an additional insight that the efficiency gains occur when the regression models are not linearly dependent across treatments. Theorem 1(b) requires correct specification, which is analogous to the requirement in Negi and Wooldridge (2020) for non-linear regression adjustment. We also show the point-wise semiparametric efficiency of the DTE estimator under correct specification in Section A.6 for the sake of completeness.
@@ 3.2 Asymptotic Distribution of DTE Estimator
(a) The log-likelihood functions $\beta\mapsto\ell_{k}(\beta;y)$ is a concave function. The link function is a canonical link function, and its inverse map $\Lambda(\cdot)$ is twice continuously differentiable with its first derivative $\lambda(\cdot)$.
@@
Suppose that Assumptions 1 and 2 hold.
(a) Then, we have, for every $k,k^{\prime}\in\mathcal{K}$ and some constant $h>0$, in $\ell^{\infty}(\mathcal{Y})$,
$\sqrt{n}(\widehat{\Delta}^{DTE}_{k,k^{\prime}}-\Delta^{DTE}_{k,k^{\prime}})\rightsquigarrow\Psi^{DTE}_{k,k^{\prime}}(\mathbb{B})\ \ \ \mathrm{and}\ \ \ \sqrt{n}(\widehat{\Delta}^{PTE}_{k,k^{\prime},h}-\Delta^{PTE}_{k,k^{\prime},h})\rightsquigarrow\Psi^{PTE}_{k,k^{\prime},h}(\mathbb{B}),$
where $\mathbb{B}$ is a mean-zero Gaussian process defined in equation (A.18) of Supplemental Material.
(b) Additionally, if the potential outcome $Y_{k}$ has a continuously differentiable distribution function with strictly positive density $f_{Y_{k}}$ on its support for every $k\in\mathcal{K}$, then
$\sqrt{n}(\widehat{\Delta}^{QTE}_{k,k^{\prime}}-\Delta^{QTE}_{k,k^{\prime}})\rightsquigarrow\psi^{DTE}_{k,k^{\prime},\gamma}(\mathbb{B}),$
where $\psi^{DTE}_{k,k^{\prime},\gamma}(h)(u):=h_{k}\circ F_{Y_{k}}^{-1}(u)/f_{Y_{k}}\circ F_{Y_{k}}^{-1}(u)-h_{k^{\prime}}\circ F_{Y_{k^{\prime}}}^{-1}(u)/f_{Y_{k^{\prime}}}\circ F_{Y_{k^{\prime}}}^{-1}(u)$ is the Hadamard derivative of $\Psi^{DTE}_{k,k^{\prime}}(\cdot)$ at $\gamma$ in direction $h:=(h_{1},\dots,h_{K})^{\top}$.
@@
Let $(S_{1},\dots,S_{n})$ be $n$ scalar, nonnegative exchangeable random variables that are independent of the original sample. The random weights $\{S_{i}\}_{i=1}^{n}$ satisfy the followings: $\mathbb{E}|S_{i}|^{2+\epsilon}<\infty$ for some $\epsilon>0$, $\bar{S}_{n}:=n^{-1}\sum_{i=1}^{n}S_{i}\to^{p}0$ and $n^{-1}\sum_{i=1}^{n}(S_{i}-\bar{S}_{n})^{2}\to^{p}1$.
@@
In the above equation, we use $\widehat{G}_{Y(k)|\bm{X}}^{\ast}$ the conditional distribution estimator based on the original sample, instead of the bootstrap sample. This improves the computational efficiency of our bootstrap inference by eliminating the need to reestimate the parameters for each bootstrap repetition. According to Theorem 2, the conditional distribution estimator has no impact on the limit distribution and we keep the estimator throughout the resampling procedure. The validity is proven in the theorem below.
@@
$\widehat{\Delta}^{DTE\ast}_{k,k^{\prime}}:=\Psi_{k,k^{\prime}}^{DTE}(\,\widehat{\gamma}^{\ast}\,),\ \ \ \widehat{\Delta}^{PTE\ast}_{k,k^{\prime},h}:=\Psi_{k,k^{\prime},h}^{PTE}(\,\widehat{\gamma}^{\ast}\,)\ \ \mathrm{and}\ \ \widehat{\Delta}^{QTE\ast}_{k,k^{\prime},h}:=\Psi_{k,k^{\prime},h}^{QTE}(\,\widehat{\gamma}^{\ast}\,).$
In the theorem provided below, we show that the exchangeable bootstrap provides a method to consistently estimate the limit process of the regression-adjusted DTE, PTE and QTE estimators, following van der Vaart and Wellner (1996).
@@
Suppose that Assumptions 1-3 hold.
(a) Then, we have, in $\ell^{\infty}(\mathcal{Y})$,
@@
(b) Additionally, if the potential outcome $Y_{k}$ has a continuously differentiable distribution function with strictly positive density $f_{Y_{k}}$ on its support for every $k\in\mathcal{K}$, then
$\sqrt{n}(\widehat{\Delta}^{QTE\ast}_{k,k^{\prime}}-\widehat{\Delta}^{QTE}_{k,k^{\prime}})\rightsquigarrow\psi^{DTE}_{k,k^{\prime},\gamma}(\mathbb{B}),$
where $\psi^{DTE}_{k,k^{\prime},\gamma}(\cdot)$ is defined in Theorem 2(b).
Theorems 2 and 3 present a methodologically rigorous and empirically useful approach for estimating DTE and PTE. Moreover, these theoretical results establish the asymptotic validity of the exchangeable bootstrap procedure for constructing confidence intervals and conducting hypothesis tests for both DTE and PTE, as well as QTE under additional conditions.
The following algorithm describes the construction of confidence intervals for regression-adjusted DTE using empirical bootstrap. The same procedure can be applied to obtain confidence intervals for regression-adjusted PTE, QTE and simple estimators.
Algorithm 1 Bootstrap confidence intervals for regression-adjusted DTE
0:   Original sample $\{(\bm{W}_{i},\bm{X}_{i},Y_{i})\}_{i=1}^{n}$
0:   $(1-\alpha)\times 100\%$ confidence intervals for the regression-adjusted DTE
0:   For each bootstrap iteration $b=1,\ldots,B$:
0:    Draw a bootstrap sample of size $n$ with replacement:  $\{(\bm{W}_{i}^{b},\bm{X}_{i}^{b},Y_{i}^{b})\}_{i=1}^{n}$ from $\{(\bm{W}_{i},\bm{X}_{i},Y_{i})\}_{i=1}^{n}$
0:    Compute regression-adjusted DTE $\widehat{\Delta}^{DTE,b}_{k,k^{\prime}}$ given the conditional  distribution estimator based on the original sample
0:   Calculate standard errors $\widehat{\Sigma}^{DTE}_{k,k^{\prime}}(y)$ using the bootstrap iterations
0:   Construct the confidence band:
$\big{\{}\widehat{\Delta}^{DTE}_{k,k^{\prime}}(y)\pm\Phi^{-1}(1-\alpha/2)\cdot\widehat{\Sigma}^{DTE}_{k,k^{\prime}}(y):y\in\mathcal{Y}\big{\}},$
 where $\Phi$ is the standard normal distribution function
In the above algorithm, the standard errors $\widehat{\Sigma}^{DTE}_{k,k^{\prime}}(y)$ can be computed by (a) using the bootstrap standard deviation of the bootstrapped DTEs $\{\widehat{\Delta}^{DTE,b}_{k,k^{\prime}}(y)\}_{b=1}^{B}$, or (b) using the rescaled bootstrap interquartile range $\big{(}q_{0.75}(y)-q_{0.25}(y)\big{)}/(z_{0.75}-z_{0.25})$, where $q_{p}(y)$ is the $p$-th quantile of the bootstrapped DTE at location $y$ and $z_{p}$ is the $p$-th quantile of the standard normal distribution.